OpenStudy (anonymous):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) Y=3x+5 X=-3y-5

OpenStudy (anonymous):

y=3(-3y-5)+5 y=-9y-15+5 10y=-10 y=-1

OpenStudy (anonymous):

x=-3.-1-5 x=3-5 x=-2 (-2,-1)

OpenStudy (anonymous):

thx nicky

OpenStudy (anonymous):

your welcome, :))